OpenStudy (anonymous):

PLEASE HELP WILL GIVE MEDAL!!! What is the value of the expression when a = 2?

OpenStudy (anonymous):

whats the equation

OpenStudy (anonymous):



OpenStudy (anonymous):

essentially its saying 5*2to the 4th power

OpenStudy (anonymous):

Is the answer 10,000?

OpenStudy (calculusxy):

\[5(2)^4 = 5 \times (2 \times 2 \times 2 \times 2)\]

OpenStudy (anonymous):

okay 2*5 is 10, 10 to the fourth power would be 10*10*10*10 = 10000

OpenStudy (anonymous):

10,000? or 80?

OpenStudy (calculusxy):

The exponent 4 only applies to 2 (which is the base).

OpenStudy (anonymous):

so 80

OpenStudy (calculusxy):

Yes :)

OpenStudy (anonymous):

thx

OpenStudy (calculusxy):

You're welcome :)

OpenStudy (anonymous):

What is the value of the expression when a = 3 and b = 2?

OpenStudy (anonymous):

i thought he was doing the easy ones

OpenStudy (calculusxy):

Where the expression?

OpenStudy (anonymous):